NOAA SATOPS Morning Report: Monday, November 05, 2012


SATOPS Morning Report: Monday, November 05, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 310

GOES
GOES-12 DOY308 Imager frame breaks at 1039Z, 1126Z and 1633Z due to RFI at WCDA.

GOES-13 DOY307 RFI AT WCDA FROM 11:39Z TO 12:16. The follow images were effected: 11:39z S. HEMI, 11:45 Full Disk, 12:15 N HEMI and 11:20z, 1146z Soundings.

All other GOES operations were nominal over the past 72 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
OPS49 DOY309 07:54 CAT III loss of 7361 s/c seconds. Problem under investigation.

All other DMSP operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
DOY308 ESPC did not receive METEOSAT-9 data November 3, 201 1645z through November 4, 2012 0215z. The outage was due to a telecom problem.

DOY308 ESPC did not receive GOES-west LRIT data from 19:00z through 19:20z

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.